# Medical Spoken Named Entity Recognition

### v2

@@ Appendix F NER Results by Entity Types
AGE BARTpho Mod. SLUE 0,70 0,79 0,74
BARTpho seqeval 0,69 0,61 0,65
BARTpho SLUE 0,69 0,78 0,73
mBART-50 Mod. SLUE 0,64 0,81 0,71
mBART-50 seqeval 0,66 0,53 0,59
mBART-50 SLUE 0,62 0,79 0,70
PhoBERT_base Mod. SLUE 0,75 0,81 0,78
PhoBERT_base seqeval 0,76 0,62 0,68
PhoBERT_base SLUE 0,74 0,80 0,77
PhoBERT_base-v2 Mod. SLUE 0,78 0,78 0,78
PhoBERT_base-v2 seqeval 0,76 0,66 0,71
PhoBERT_base-v2 SLUE 0,77 0,77 0,77
PhoBERT_large Mod. SLUE 0,77 0,81 0,79
PhoBERT_large seqeval 0,77 0,66 0,71
PhoBERT_large SLUE 0,76 0,80 0,78
ViDeBERTa_base Mod. SLUE 0,59 0,05 0,09
ViDeBERTa_base seqeval 0,03 0,29 0,05
ViDeBERTa_base SLUE 0,50 0,04 0,07
ViT5_base Mod. SLUE 0,71 0,79 0,75
ViT5_base seqeval 0,73 0,63 0,68
ViT5_base SLUE 0,69 0,77 0,73
XLM-R_base Mod. SLUE 0,69 0,79 0,73
XLM-R_base seqeval 0,71 0,56 0,63
XLM-R_base SLUE 0,68 0,77 0,72
XLM-R_large Mod. SLUE 0,78 0,77 0,78
XLM-R_large seqeval 0,75 0,67 0,71
XLM-R_large SLUE 0,78 0,77 0,77
DATETIME BARTpho Mod. SLUE 0,76 0,76 0,76
BARTpho seqeval 0,62 0,66 0,64
BARTpho SLUE 0,74 0,75 0,75
mBART-50 Mod. SLUE 0,83 0,74 0,78
mBART-50 seqeval 0,67 0,75 0,71
mBART-50 SLUE 0,82 0,73 0,77
PhoBERT_base Mod. SLUE 0,80 0,83 0,82
PhoBERT_base seqeval 0,77 0,70 0,74
PhoBERT_base SLUE 0,80 0,82 0,81
PhoBERT_base-v2 Mod. SLUE 0,81 0,84 0,83
PhoBERT_base-v2 seqeval 0,78 0,73 0,75
PhoBERT_base-v2 SLUE 0,81 0,83 0,82
PhoBERT_large Mod. SLUE 0,83 0,82 0,83
PhoBERT_large seqeval 0,76 0,75 0,76
PhoBERT_large SLUE 0,83 0,81 0,82
ViDeBERTa_base Mod. SLUE 0,62 0,68 0,65
ViDeBERTa_base seqeval 0,58 0,43 0,50
ViDeBERTa_base SLUE 0,60 0,65 0,62
ViT5_base Mod. SLUE 0,74 0,77 0,75
ViT5_base seqeval 0,69 0,67 0,68
ViT5_base SLUE 0,72 0,75 0,74
XLM-R_base Mod. SLUE 0,81 0,78 0,80
XLM-R_base seqeval 0,72 0,70 0,71
XLM-R_base SLUE 0,80 0,77 0,78
XLM-R_large Mod. SLUE 0,85 0,82 0,83
XLM-R_large seqeval 0,76 0,77 0,77
XLM-R_large SLUE 0,84 0,81 0,82
DIAGNOSTICS BARTpho Mod. SLUE 0,66 0,82 0,73
BARTpho seqeval 0,72 0,65 0,68
BARTpho SLUE 0,65 0,82 0,73
mBART-50 Mod. SLUE 0,58 0,81 0,68
mBART-50 seqeval 0,71 0,59 0,65
mBART-50 SLUE 0,57 0,80 0,66
PhoBERT_base Mod. SLUE 0,66 0,81 0,73
PhoBERT_base seqeval 0,75 0,64 0,69
PhoBERT_base SLUE 0,66 0,80 0,72
PhoBERT_base-v2 Mod. SLUE 0,66 0,82 0,73
PhoBERT_base-v2 seqeval 0,77 0,65 0,70
PhoBERT_base-v2 SLUE 0,66 0,82 0,73
PhoBERT_large Mod. SLUE 0,70 0,83 0,76
PhoBERT_large seqeval 0,78 0,69 0,73
PhoBERT_large SLUE 0,69 0,83 0,75
ViDeBERTa_base Mod. SLUE 0,53 0,60 0,56
ViDeBERTa_base seqeval 0,55 0,39 0,45
ViDeBERTa_base SLUE 0,51 0,58 0,54
ViT5_base Mod. SLUE 0,60 0,80 0,69
ViT5_base seqeval 0,71 0,62 0,67
ViT5_base SLUE 0,59 0,78 0,67
XLM-R_base Mod. SLUE 0,61 0,82 0,70
XLM-R_base seqeval 0,75 0,58 0,65
XLM-R_base SLUE 0,60 0,82 0,69
XLM-R_large Mod. SLUE 0,69 0,78 0,73
XLM-R_large seqeval 0,78 0,69 0,73
XLM-R_large SLUE 0,69 0,78 0,73
DISEASESYMTOM BARTpho Mod. SLUE 0,75 0,81 0,78
BARTpho seqeval 0,63 0,60 0,61
BARTpho SLUE 0,73 0,78 0,75
mBART-50 Mod. SLUE 0,71 0,73 0,72
mBART-50 seqeval 0,57 0,56 0,57
mBART-50 SLUE 0,68 0,69 0,69
PhoBERT_base Mod. SLUE 0,79 0,83 0,81
PhoBERT_base seqeval 0,70 0,60 0,65
PhoBERT_base SLUE 0,77 0,80 0,78
PhoBERT_base-v2 Mod. SLUE 0,79 0,85 0,82
PhoBERT_base-v2 seqeval 0,73 0,61 0,66
PhoBERT_base-v2 SLUE 0,77 0,82 0,79
PhoBERT_large Mod. SLUE 0,82 0,81 0,82
PhoBERT_large seqeval 0,71 0,64 0,67
PhoBERT_large SLUE 0,79 0,79 0,79
ViDeBERTa_base Mod. SLUE 0,68 0,52 0,59
ViDeBERTa_base seqeval 0,35 0,38 0,37
ViDeBERTa_base SLUE 0,62 0,47 0,54
ViT5_base Mod. SLUE 0,78 0,84 0,81
ViT5_base seqeval 0,73 0,67 0,70
ViT5_base SLUE 0,77 0,82 0,79
XLM-R_base Mod. SLUE 0,75 0,83 0,79
XLM-R_base seqeval 0,69 0,56 0,61
XLM-R_base SLUE 0,72 0,79 0,75
XLM-R_large Mod. SLUE 0,83 0,81 0,82
XLM-R_large seqeval 0,70 0,66 0,68
XLM-R_large SLUE 0,81 0,79 0,80
DRUGCHEMICAL BARTpho Mod. SLUE 0,73 0,82 0,77
BARTpho seqeval 0,75 0,70 0,73
BARTpho SLUE 0,73 0,81 0,77
mBART-50 Mod. SLUE 0,79 0,69 0,74
mBART-50 seqeval 0,71 0,76 0,73
mBART-50 SLUE 0,79 0,69 0,74
PhoBERT_base Mod. SLUE 0,80 0,93 0,86
PhoBERT_base seqeval 0,91 0,77 0,83
PhoBERT_base SLUE 0,79 0,93 0,86
PhoBERT_base-v2 Mod. SLUE 0,83 0,93 0,88
PhoBERT_base-v2 seqeval 0,91 0,80 0,85
PhoBERT_base-v2 SLUE 0,83 0,93 0,88
PhoBERT_large Mod. SLUE 0,74 0,93 0,82
PhoBERT_large seqeval 0,93 0,72 0,81
PhoBERT_large SLUE 0,74 0,93 0,82
ViDeBERTa_base Mod. SLUE 0,65 0,41 0,51
ViDeBERTa_base seqeval 0,31 0,41 0,35
ViDeBERTa_base SLUE 0,60 0,38 0,47
ViT5_base Mod. SLUE 0,75 0,87 0,80
ViT5_base seqeval 0,83 0,74 0,78
ViT5_base SLUE 0,75 0,86 0,80
XLM-R_base Mod. SLUE 0,81 0,74 0,77
XLM-R_base seqeval 0,76 0,76 0,76
XLM-R_base SLUE 0,80 0,74 0,77
XLM-R_large Mod. SLUE 0,85 0,92 0,88
XLM-R_large seqeval 0,91 0,79 0,85
XLM-R_large SLUE 0,85 0,92 0,88
FOODDRINK BARTpho Mod. SLUE 0,67 0,83 0,74
BARTpho seqeval 0,69 0,59 0,64
BARTpho SLUE 0,67 0,83 0,74
mBART-50 Mod. SLUE 0,56 0,67 0,61
mBART-50 seqeval 0,57 0,52 0,54
mBART-50 SLUE 0,56 0,67 0,61
PhoBERT_base Mod. SLUE 0,78 0,84 0,81
PhoBERT_base seqeval 0,74 0,67 0,70
PhoBERT_base SLUE 0,78 0,84 0,81
PhoBERT_base-v2 Mod. SLUE 0,82 0,89 0,85
PhoBERT_base-v2 seqeval 0,83 0,76 0,79
PhoBERT_base-v2 SLUE 0,82 0,89 0,85
PhoBERT_large Mod. SLUE 0,80 0,91 0,85
PhoBERT_large seqeval 0,85 0,75 0,80
PhoBERT_large SLUE 0,80 0,91 0,85
ViDeBERTa_base Mod. SLUE 0,22 0,38 0,28
ViDeBERTa_base seqeval 0,24 0,12 0,16
ViDeBERTa_base SLUE 0,19 0,32 0,23
ViT5_base Mod. SLUE 0,70 0,86 0,77
ViT5_base seqeval 0,75 0,66 0,70
ViT5_base SLUE 0,70 0,86 0,77
XLM-R_base Mod. SLUE 0,64 0,88 0,74
XLM-R_base seqeval 0,82 0,58 0,68
XLM-R_base SLUE 0,62 0,85 0,72
XLM-R_large Mod. SLUE 0,88 0,81 0,84
XLM-R_large seqeval 0,75 0,83 0,79
XLM-R_large SLUE 0,88 0,81 0,84
GENDER BARTpho Mod. SLUE 0,83 0,90 0,86
BARTpho seqeval 0,85 0,78 0,81
BARTpho SLUE 0,82 0,90 0,86
mBART-50 Mod. SLUE 0,83 0,87 0,85
mBART-50 seqeval 0,77 0,75 0,76
mBART-50 SLUE 0,82 0,86 0,84
PhoBERT_base Mod. SLUE 0,83 0,90 0,86
PhoBERT_base seqeval 0,76 0,74 0,75
PhoBERT_base SLUE 0,83 0,90 0,86
PhoBERT_base-v2 Mod. SLUE 0,84 0,89 0,86
PhoBERT_base-v2 seqeval 0,83 0,77 0,80
PhoBERT_base-v2 SLUE 0,84 0,89 0,86
PhoBERT_large Mod. SLUE 0,83 0,90 0,87
PhoBERT_large seqeval 0,87 0,79 0,83
PhoBERT_large SLUE 0,83 0,90 0,86
ViDeBERTa_base Mod. SLUE 0,00 0,00 0,00
ViDeBERTa_base seqeval 0,00 0,00 0,00
ViDeBERTa_base SLUE 0,00 0,00 0,00
ViT5_base Mod. SLUE 0,81 0,82 0,82
ViT5_base seqeval 0,67 0,71 0,69
ViT5_base SLUE 0,81 0,82 0,81
XLM-R_base Mod. SLUE 0,79 0,72 0,75
XLM-R_base seqeval 0,56 0,49 0,52
XLM-R_base SLUE 0,78 0,71 0,75
XLM-R_large Mod. SLUE 0,91 0,90 0,90
XLM-R_large seqeval 0,82 0,79 0,80
XLM-R_large SLUE 0,91 0,89 0,90
LOCATION BARTpho Mod. SLUE 0,77 0,88 0,82
BARTpho seqeval 0,75 0,68 0,71
BARTpho SLUE 0,76 0,87 0,81
mBART-50 Mod. SLUE 0,74 0,83 0,78
mBART-50 seqeval 0,73 0,68 0,71
mBART-50 SLUE 0,74 0,82 0,78
PhoBERT_base Mod. SLUE 0,75 0,92 0,82
PhoBERT_base seqeval 0,81 0,63 0,71
PhoBERT_base SLUE 0,74 0,91 0,81
PhoBERT_base-v2 Mod. SLUE 0,78 0,95 0,86
PhoBERT_base-v2 seqeval 0,86 0,69 0,77
PhoBERT_base-v2 SLUE 0,78 0,94 0,86
PhoBERT_large Mod. SLUE 0,80 0,91 0,85
PhoBERT_large seqeval 0,82 0,69 0,75
PhoBERT_large SLUE 0,80 0,90 0,84
ViDeBERTa_base Mod. SLUE 0,78 0,33 0,47
ViDeBERTa_base seqeval 0,32 0,60 0,42
ViDeBERTa_base SLUE 0,77 0,33 0,46
ViT5_base Mod. SLUE 0,74 0,92 0,82
ViT5_base seqeval 0,83 0,65 0,73
ViT5_base SLUE 0,73 0,92 0,81
XLM-R_base Mod. SLUE 0,75 0,84 0,79
XLM-R_base seqeval 0,75 0,59 0,66
XLM-R_base SLUE 0,74 0,82 0,78
XLM-R_large Mod. SLUE 0,85 0,93 0,89
XLM-R_large seqeval 0,87 0,75 0,81
XLM-R_large SLUE 0,85 0,93 0,88
MEDDEVICETECHNIQUE BARTpho Mod. SLUE 0,54 0,21 0,30
BARTpho seqeval 0,16 0,39 0,23
BARTpho SLUE 0,48 0,18 0,26
mBART-50 Mod. SLUE 0,45 0,09 0,15
mBART-50 seqeval 0,08 0,35 0,13
mBART-50 SLUE 0,36 0,07 0,12
PhoBERT_base Mod. SLUE 0,55 0,38 0,45
PhoBERT_base seqeval 0,24 0,27 0,26
PhoBERT_base SLUE 0,49 0,34 0,40
PhoBERT_base-v2 Mod. SLUE 0,59 0,38 0,46
PhoBERT_base-v2 seqeval 0,27 0,35 0,31
PhoBERT_base-v2 SLUE 0,53 0,34 0,41
PhoBERT_large Mod. SLUE 0,61 0,30 0,40
PhoBERT_large seqeval 0,22 0,35 0,27
PhoBERT_large SLUE 0,55 0,27 0,36
ViDeBERTa_base Mod. SLUE 0,34 0,14 0,20
ViDeBERTa_base seqeval 0,06 0,09 0,07
ViDeBERTa_base SLUE 0,24 0,10 0,14
ViT5_base Mod. SLUE 0,52 0,21 0,30
ViT5_base seqeval 0,16 0,37 0,23
ViT5_base SLUE 0,47 0,19 0,27
XLM-R_base Mod. SLUE 0,48 0,27 0,34
XLM-R_base seqeval 0,12 0,18 0,14
XLM-R_base SLUE 0,41 0,23 0,29
XLM-R_large Mod. SLUE 0,58 0,28 0,38
XLM-R_large seqeval 0,20 0,32 0,25
XLM-R_large SLUE 0,51 0,25 0,33
OCCUPATION BARTpho Mod. SLUE 0,91 0,93 0,92
BARTpho seqeval 0,87 0,87 0,87
BARTpho SLUE 0,91 0,92 0,92
mBART-50 Mod. SLUE 0,97 0,93 0,95
mBART-50 seqeval 0,91 0,95 0,93
mBART-50 SLUE 0,97 0,93 0,95
PhoBERT_base Mod. SLUE 0,95 0,96 0,95
PhoBERT_base seqeval 0,94 0,92 0,93
PhoBERT_base SLUE 0,95 0,96 0,95
PhoBERT_base-v2 Mod. SLUE 0,96 0,96 0,96
PhoBERT_base-v2 seqeval 0,95 0,93 0,94
PhoBERT_base-v2 SLUE 0,96 0,96 0,96
PhoBERT_large Mod. SLUE 0,97 0,95 0,96
PhoBERT_large seqeval 0,93 0,94 0,94
PhoBERT_large SLUE 0,97 0,95 0,96
ViDeBERTa_base Mod. SLUE 0,96 0,82 0,89
ViDeBERTa_base seqeval 0,81 0,91 0,86
ViDeBERTa_base SLUE 0,96 0,81 0,88
ViT5_base Mod. SLUE 0,95 0,93 0,94
ViT5_base seqeval 0,92 0,94 0,93
ViT5_base SLUE 0,95 0,93 0,94
XLM-R_base Mod. SLUE 0,88 0,96 0,92
XLM-R_base seqeval 0,93 0,83 0,88
[…]

### v3

@@ Appendix E NER Results by Entity Types
AGE BARTpho Mod. SLUE 0.70 0.79 0.74
BARTpho seqeval 0.69 0.61 0.65
BARTpho SLUE 0.69 0.78 0.73
mBART-50 Mod. SLUE 0.64 0.81 0.71
mBART-50 seqeval 0.66 0.53 0.59
mBART-50 SLUE 0.62 0.79 0.70
PhoBERT_base Mod. SLUE 0.75 0.81 0.78
PhoBERT_base seqeval 0.76 0.62 0.68
PhoBERT_base SLUE 0.74 0.80 0.77
PhoBERT_base-v2 Mod. SLUE 0.78 0.78 0.78
PhoBERT_base-v2 seqeval 0.76 0.66 0.71
PhoBERT_base-v2 SLUE 0.77 0.77 0.77
PhoBERT_large Mod. SLUE 0.77 0.81 0.79
PhoBERT_large seqeval 0.77 0.66 0.71
PhoBERT_large SLUE 0.76 0.80 0.78
ViDeBERTa_base Mod. SLUE 0.59 0.05 0.09
ViDeBERTa_base seqeval 0.03 0.29 0.05
ViDeBERTa_base SLUE 0.50 0.04 0.07
ViT5_base Mod. SLUE 0.71 0.79 0.75
ViT5_base seqeval 0.73 0.63 0.68
ViT5_base SLUE 0.69 0.77 0.73
XLM-R_base Mod. SLUE 0.69 0.79 0.73
XLM-R_base seqeval 0.71 0.56 0.63
XLM-R_base SLUE 0.68 0.77 0.72
XLM-R_large Mod. SLUE 0.78 0.77 0.78
XLM-R_large seqeval 0.75 0.67 0.71
XLM-R_large SLUE 0.78 0.77 0.77
Table 5: NER results of AGE entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
DATETIME BARTpho Mod. SLUE 0.76 0.76 0.76
BARTpho seqeval 0.62 0.66 0.64
BARTpho SLUE 0.74 0.75 0.75
mBART-50 Mod. SLUE 0.83 0.74 0.78
mBART-50 seqeval 0.67 0.75 0.71
mBART-50 SLUE 0.82 0.73 0.77
PhoBERT_base Mod. SLUE 0.80 0.83 0.82
PhoBERT_base seqeval 0.77 0.70 0.74
PhoBERT_base SLUE 0.80 0.82 0.81
PhoBERT_base-v2 Mod. SLUE 0.81 0.84 0.83
PhoBERT_base-v2 seqeval 0.78 0.73 0.75
PhoBERT_base-v2 SLUE 0.81 0.83 0.82
PhoBERT_large Mod. SLUE 0.83 0.82 0.83
PhoBERT_large seqeval 0.76 0.75 0.76
PhoBERT_large SLUE 0.83 0.81 0.82
ViDeBERTa_base Mod. SLUE 0.62 0.68 0.65
ViDeBERTa_base seqeval 0.58 0.43 0.50
ViDeBERTa_base SLUE 0.60 0.65 0.62
ViT5_base Mod. SLUE 0.74 0.77 0.75
ViT5_base seqeval 0.69 0.67 0.68
ViT5_base SLUE 0.72 0.75 0.74
XLM-R_base Mod. SLUE 0.81 0.78 0.80
XLM-R_base seqeval 0.72 0.70 0.71
XLM-R_base SLUE 0.80 0.77 0.78
XLM-R_large Mod. SLUE 0.85 0.82 0.83
XLM-R_large seqeval 0.76 0.77 0.77
XLM-R_large SLUE 0.84 0.81 0.82
Table 6: NER results of DATETIME entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
DIAGNOSTICS BARTpho Mod. SLUE 0.66 0.82 0.73
BARTpho seqeval 0.72 0.65 0.68
BARTpho SLUE 0.65 0.82 0.73
mBART-50 Mod. SLUE 0.58 0.81 0.68
mBART-50 seqeval 0.71 0.59 0.65
mBART-50 SLUE 0.57 0.80 0.66
PhoBERT_base Mod. SLUE 0.66 0.81 0.73
PhoBERT_base seqeval 0.75 0.64 0.69
PhoBERT_base SLUE 0.66 0.80 0.72
PhoBERT_base-v2 Mod. SLUE 0.66 0.82 0.73
PhoBERT_base-v2 seqeval 0.77 0.65 0.70
PhoBERT_base-v2 SLUE 0.66 0.82 0.73
PhoBERT_large Mod. SLUE 0.70 0.83 0.76
PhoBERT_large seqeval 0.78 0.69 0.73
PhoBERT_large SLUE 0.69 0.83 0.75
ViDeBERTa_base Mod. SLUE 0.53 0.60 0.56
ViDeBERTa_base seqeval 0.55 0.39 0.45
ViDeBERTa_base SLUE 0.51 0.58 0.54
ViT5_base Mod. SLUE 0.60 0.80 0.69
ViT5_base seqeval 0.71 0.62 0.67
ViT5_base SLUE 0.59 0.78 0.67
XLM-R_base Mod. SLUE 0.61 0.82 0.70
XLM-R_base seqeval 0.75 0.58 0.65
XLM-R_base SLUE 0.60 0.82 0.69
XLM-R_large Mod. SLUE 0.69 0.78 0.73
XLM-R_large seqeval 0.78 0.69 0.73
XLM-R_large SLUE 0.69 0.78 0.73
Table 7: NER results of DIAGNOSTICS entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
DISEASESYMTOM BARTpho Mod. SLUE 0.75 0.81 0.78
BARTpho seqeval 0.63 0.60 0.61
BARTpho SLUE 0.73 0.78 0.75
mBART-50 Mod. SLUE 0.71 0.73 0.72
mBART-50 seqeval 0.57 0.56 0.57
mBART-50 SLUE 0.68 0.69 0.69
PhoBERT_base Mod. SLUE 0.79 0.83 0.81
PhoBERT_base seqeval 0.70 0.60 0.65
PhoBERT_base SLUE 0.77 0.80 0.78
PhoBERT_base-v2 Mod. SLUE 0.79 0.85 0.82
PhoBERT_base-v2 seqeval 0.73 0.61 0.66
PhoBERT_base-v2 SLUE 0.77 0.82 0.79
PhoBERT_large Mod. SLUE 0.82 0.81 0.82
PhoBERT_large seqeval 0.71 0.64 0.67
PhoBERT_large SLUE 0.79 0.79 0.79
ViDeBERTa_base Mod. SLUE 0.68 0.52 0.59
ViDeBERTa_base seqeval 0.35 0.38 0.37
ViDeBERTa_base SLUE 0.62 0.47 0.54
ViT5_base Mod. SLUE 0.78 0.84 0.81
ViT5_base seqeval 0.73 0.67 0.70
ViT5_base SLUE 0.77 0.82 0.79
XLM-R_base Mod. SLUE 0.75 0.83 0.79
XLM-R_base seqeval 0.69 0.56 0.61
XLM-R_base SLUE 0.72 0.79 0.75
XLM-R_large Mod. SLUE 0.83 0.81 0.82
XLM-R_large seqeval 0.70 0.66 0.68
XLM-R_large SLUE 0.81 0.79 0.80
Table 8: NER results of DISEASESYMTOM entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
DRUGCHEMICAL BARTpho Mod. SLUE 0.73 0.82 0.77
BARTpho seqeval 0.75 0.70 0.73
BARTpho SLUE 0.73 0.81 0.77
mBART-50 Mod. SLUE 0.79 0.69 0.74
mBART-50 seqeval 0.71 0.76 0.73
mBART-50 SLUE 0.79 0.69 0.74
PhoBERT_base Mod. SLUE 0.80 0.93 0.86
PhoBERT_base seqeval 0.91 0.77 0.83
PhoBERT_base SLUE 0.79 0.93 0.86
PhoBERT_base-v2 Mod. SLUE 0.83 0.93 0.88
PhoBERT_base-v2 seqeval 0.91 0.80 0.85
PhoBERT_base-v2 SLUE 0.83 0.93 0.88
PhoBERT_large Mod. SLUE 0.74 0.93 0.82
PhoBERT_large seqeval 0.93 0.72 0.81
PhoBERT_large SLUE 0.74 0.93 0.82
ViDeBERTa_base Mod. SLUE 0.65 0.41 0.51
ViDeBERTa_base seqeval 0.31 0.41 0.35
ViDeBERTa_base SLUE 0.60 0.38 0.47
ViT5_base Mod. SLUE 0.75 0.87 0.80
ViT5_base seqeval 0.83 0.74 0.78
ViT5_base SLUE 0.75 0.86 0.80
XLM-R_base Mod. SLUE 0.81 0.74 0.77
XLM-R_base seqeval 0.76 0.76 0.76
XLM-R_base SLUE 0.80 0.74 0.77
XLM-R_large Mod. SLUE 0.85 0.92 0.88
XLM-R_large seqeval 0.91 0.79 0.85
XLM-R_large SLUE 0.85 0.92 0.88
Table 9: NER results of DRUGCHEMICAL entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
FOODDRINK BARTpho Mod. SLUE 0.67 0.83 0.74
BARTpho seqeval 0.69 0.59 0.64
BARTpho SLUE 0.67 0.83 0.74
mBART-50 Mod. SLUE 0.56 0.67 0.61
mBART-50 seqeval 0.57 0.52 0.54
mBART-50 SLUE 0.56 0.67 0.61
PhoBERT_base Mod. SLUE 0.78 0.84 0.81
PhoBERT_base seqeval 0.74 0.67 0.70
PhoBERT_base SLUE 0.78 0.84 0.81
PhoBERT_base-v2 Mod. SLUE 0.82 0.89 0.85
PhoBERT_base-v2 seqeval 0.83 0.76 0.79
PhoBERT_base-v2 SLUE 0.82 0.89 0.85
PhoBERT_large Mod. SLUE 0.80 0.91 0.85
PhoBERT_large seqeval 0.85 0.75 0.80
PhoBERT_large SLUE 0.80 0.91 0.85
ViDeBERTa_base Mod. SLUE 0.22 0.38 0.28
ViDeBERTa_base seqeval 0.24 0.12 0.16
ViDeBERTa_base SLUE 0.19 0.32 0.23
ViT5_base Mod. SLUE 0.70 0.86 0.77
ViT5_base seqeval 0.75 0.66 0.70
ViT5_base SLUE 0.70 0.86 0.77
XLM-R_base Mod. SLUE 0.64 0.88 0.74
XLM-R_base seqeval 0.82 0.58 0.68
XLM-R_base SLUE 0.62 0.85 0.72
XLM-R_large Mod. SLUE 0.88 0.81 0.84
XLM-R_large seqeval 0.75 0.83 0.79
XLM-R_large SLUE 0.88 0.81 0.84
Table 10: NER results of FOODDRINK entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
GENDER BARTpho Mod. SLUE 0.83 0.90 0.86
BARTpho seqeval 0.85 0.78 0.81
BARTpho SLUE 0.82 0.90 0.86
mBART-50 Mod. SLUE 0.83 0.87 0.85
mBART-50 seqeval 0.77 0.75 0.76
mBART-50 SLUE 0.82 0.86 0.84
PhoBERT_base Mod. SLUE 0.83 0.90 0.86
PhoBERT_base seqeval 0.76 0.74 0.75
PhoBERT_base SLUE 0.83 0.90 0.86
PhoBERT_base-v2 Mod. SLUE 0.84 0.89 0.86
PhoBERT_base-v2 seqeval 0.83 0.77 0.80
PhoBERT_base-v2 SLUE 0.84 0.89 0.86
PhoBERT_large Mod. SLUE 0.83 0.90 0.87
PhoBERT_large seqeval 0.87 0.79 0.83
PhoBERT_large SLUE 0.83 0.90 0.86
ViDeBERTa_base Mod. SLUE 0.00 0.00 0.00
ViDeBERTa_base seqeval 0.00 0.00 0.00
ViDeBERTa_base SLUE 0.00 0.00 0.00
ViT5_base Mod. SLUE 0.81 0.82 0.82
ViT5_base seqeval 0.67 0.71 0.69
ViT5_base SLUE 0.81 0.82 0.81
XLM-R_base Mod. SLUE 0.79 0.72 0.75
XLM-R_base seqeval 0.56 0.49 0.52
XLM-R_base SLUE 0.78 0.71 0.75
XLM-R_large Mod. SLUE 0.91 0.90 0.90
XLM-R_large seqeval 0.82 0.79 0.80
XLM-R_large SLUE 0.91 0.89 0.90
Table 11: NER results of GENDER entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
LOCATION BARTpho Mod. SLUE 0.77 0.88 0.82
BARTpho seqeval 0.75 0.68 0.71
BARTpho SLUE 0.76 0.87 0.81
mBART-50 Mod. SLUE 0.74 0.83 0.78
mBART-50 seqeval 0.73 0.68 0.71
mBART-50 SLUE 0.74 0.82 0.78
PhoBERT_base Mod. SLUE 0.75 0.92 0.82
PhoBERT_base seqeval 0.81 0.63 0.71
PhoBERT_base SLUE 0.74 0.91 0.81
PhoBERT_base-v2 Mod. SLUE 0.78 0.95 0.86
PhoBERT_base-v2 seqeval 0.86 0.69 0.77
PhoBERT_base-v2 SLUE 0.78 0.94 0.86
PhoBERT_large Mod. SLUE 0.80 0.91 0.85
PhoBERT_large seqeval 0.82 0.69 0.75
PhoBERT_large SLUE 0.80 0.90 0.84
ViDeBERTa_base Mod. SLUE 0.78 0.33 0.47
ViDeBERTa_base seqeval 0.32 0.60 0.42
ViDeBERTa_base SLUE 0.77 0.33 0.46
ViT5_base Mod. SLUE 0.74 0.92 0.82
ViT5_base seqeval 0.83 0.65 0.73
ViT5_base SLUE 0.73 0.92 0.81
XLM-R_base Mod. SLUE 0.75 0.84 0.79
XLM-R_base seqeval 0.75 0.59 0.66
XLM-R_base SLUE 0.74 0.82 0.78
XLM-R_large Mod. SLUE 0.85 0.93 0.89
XLM-R_large seqeval 0.87 0.75 0.81
XLM-R_large SLUE 0.85 0.93 0.88
Table 12: NER results of LOCATION entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
LOCATION BARTpho Mod. SLUE 0.54 0.21 0.30
BARTpho seqeval 0.16 0.39 0.23
BARTpho SLUE 0.48 0.18 0.26
mBART-50 Mod. SLUE 0.45 0.09 0.15
mBART-50 seqeval 0.08 0.35 0.13
mBART-50 SLUE 0.36 0.07 0.12
PhoBERT_base Mod. SLUE 0.55 0.38 0.45
PhoBERT_base seqeval 0.24 0.27 0.26
PhoBERT_base SLUE 0.49 0.34 0.40
PhoBERT_base-v2 Mod. SLUE 0.59 0.38 0.46
PhoBERT_base-v2 seqeval 0.27 0.35 0.31
PhoBERT_base-v2 SLUE 0.53 0.34 0.41
PhoBERT_large Mod. SLUE 0.61 0.30 0.40
PhoBERT_large seqeval 0.22 0.35 0.27
PhoBERT_large SLUE 0.55 0.27 0.36
ViDeBERTa_base Mod. SLUE 0.34 0.14 0.20
ViDeBERTa_base seqeval 0.06 0.09 0.07
ViDeBERTa_base SLUE 0.24 0.10 0.14
ViT5_base Mod. SLUE 0.52 0.21 0.30
ViT5_base seqeval 0.16 0.37 0.23
ViT5_base SLUE 0.47 0.19 0.27
XLM-R_base Mod. SLUE 0.48 0.27 0.34
XLM-R_base seqeval 0.12 0.18 0.14
XLM-R_base SLUE 0.41 0.23 0.29
XLM-R_large Mod. SLUE 0.58 0.28 0.38
XLM-R_large seqeval 0.20 0.32 0.25
XLM-R_large SLUE 0.51 0.25 0.33
Table 13: NER results of MEDDEVICETECHNIQUE entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
OCCUPATION BARTpho Mod. SLUE 0.91 0.93 0.92
BARTpho seqeval 0.87 0.87 0.87
BARTpho SLUE 0.91 0.92 0.92
mBART-50 Mod. SLUE 0.97 0.93 0.95
mBART-50 seqeval 0.91 0.95 0.93
mBART-50 SLUE 0.97 0.93 0.95
PhoBERT_base Mod. SLUE 0.95 0.96 0.95
PhoBERT_base seqeval 0.94 0.92 0.93
PhoBERT_base SLUE 0.95 0.96 0.95
PhoBERT_base-v2 Mod. SLUE 0.96 0.96 0.96
PhoBERT_base-v2 seqeval 0.95 0.93 0.94
PhoBERT_base-v2 SLUE 0.96 0.96 0.96
PhoBERT_large Mod. SLUE 0.97 0.95 0.96
[…]